Find the centroid of the area bounded by

y2 = x3, y = 2x


1
Expert's answer
2021-05-07T10:07:05-0400

First, we find the points of intersection of the graphs "y^2=x^3" and y=2x.

"y^2=x^3=(2x)^2=4x^2" : x=0, y=0 and x=4, y=8.

On the segment "0\leq x\leq 4" we have "2x\geq x^{3/2}".

Let "\Omega" denotes the region "\{x^{3/2}\leq y\leq 2x, 0\leq x\leq 4\}".

The centroid "C" of "\Omega" has coordinates "x_C=I_x/A,\ y_C=I_y/A", where "I_x=\int\int_{\Omega}xdxdy", "I_y=\int\int_{\Omega}ydxdy", "A=\int\int_{\Omega}dxdy".

Calculate these integrals:

"A=\int\int_{\Omega}dxdy=\int\limits_0^4\left(\int\limits_{x^{3/2}}^{2x}dy\right)dx=\int\limits_0^4(2x-x^{3/2})dx=(x^2-2/5x^{5/2})|_0^4 = 16-12.8=3.2"

"I_x=\int\int_{\Omega}xdxdy=\int\limits_0^4\left(\int\limits_{x^{3/2}}^{2x}dy\right)xdx=\int\limits_0^4x(2x-x^{3/2})dx=(\frac{2}{3}x^3-\frac{2}{7}x^{7/2})|_0^4=128/3-256/7=128/21"

"I_y=\int\int_{\Omega}ydxdy=\int\limits_0^4\left(\int\limits_{x^{3/2}}^{2x}ydy\right)dx=\int\limits_0^4\frac{y^2}{2}\left|^{2x}_{x^{3/2}}\right.dx=\int\limits_0^4(2x^2-x^3/2)dx=\left(\frac{2}{3}x^3-\frac{1}{8}x^4\right)\left|_0^4\right.=128/3-32=32/3"

Now we find the coordinates of the centroid:

"x_C=(128\/21)\/3.2=40\/21"

"y_C=(32\/3)\/3.2=10\/3"


Answer. "C=\left(\frac{40}{21},\frac{10}{3}\right)".